OpenStudy (anonymous):

CHECK THIS PLEASE!! write using radical notation (64x^3)^(-1/3)

OpenStudy (anonymous):

\[\sqrt[3]{-64x^3}\]

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

actually u dont write it as negative 64. \[(64^3^-1/3 = (64)^ 3 . -1/3 => 64^-1 = 1/64.\]

OpenStudy (anonymous):

im not trying to simplify it I'm just to put it into radical notation form

OpenStudy (anonymous):

well nt sure then :P

jimthompson5910 (jim_thompson5910):

\[\large (64x^3)^{-1/3}\] \[\large \frac{1}{(64x^3)^{1/3}}\] \[\large \frac{1}{\sqrt[3]{64x^3}}\] You can go further to simplify, but I think this is where they want you to stop

OpenStudy (anonymous):

thank you very much

jimthompson5910 (jim_thompson5910):

np